OpenStudy (anonymous):

Create your own quadratic equation and demonstrate how it would be solved by graphing, factoring, the quadratic formula, and by completing the square.

OpenStudy (anonymous):

x^2-6x+9 = 0 (x-3)^2 = 0 ----> x = 3 x^2-6x = -9 ----> (x-3)^2 - 9 = -9 ---> (x+3)^2 = 0 (6+-sqrt(6^2-4*9))/2 = 6/2 = 3

OpenStudy (anonymous):

u just copy and pasted that from yahoo answers